# A Review On Safe Reinforcement Learning Using Lyapunov and Barrier Functions

### v2

@@ 5 Discussion
Safe RL with Lyapunov or barrier functions is an effective and promising field that aims to ensure the safety of an agent while it learns a task through RL. Table 12 and 3 summarize the methodology of using Lyapunov and barrier functions in RL.
@@ 5.1 Simulation Benchmarks
Simulated benchmarks have proven to be a norm in implementing and validating data-driven ML and RL algorithms over the last few decades[127]. The key challenges involved in creating a successful benchmark for these data-driven algorithm are reproducibility, ease of implementation, computational complexity (should be less to allow testing on all scales of hardware), incorporating stochastic behavior and sufficient environment complexity (indicating the environment goals should be sufficiently complex). We review some of the existing environments for benchmarking safe RL and address its usefulness in validating methodologies incorporating Lyapunov and barrier functions.
(a) (A) Island Navigation environment [128]
(b) (B) Constraints in Safe Gym [129]
@@
The earliest benchmark for safe RL approached the problem as a safe exploration towards a goal. DeepMind introduced AI Safety Gridworlds [128] and OpenAI released Safety Gym [129] as benchmarks for safe exploration problem in RL. Safety Gym proposes a constrained RL formulation for safe exploration and provide gym [130] environments [131] for high-dimensional continuous control in safe RL. Safety Gym was upgraded by introducing Safety-Gymnasium [132] with increased functionality by using MuJoCo physics engine [133], higher customization, extended single and multi-agent scenarios. Safety-Gymnasium is built on Gymnasium [134] RL environment structure and allows for easy integration with RL libraries. Safe-Gymnasium also proposes SafePO which provides single and multi-agent approaches for constrained safe RL problems. The ideal benchmarking environment for reviewed methodologies should incorporate constraints, external disturbances, nonlinear dynamical models and low-level oriented tasks. For this reason Safe-Control-Gym [135] appears as the most suitable benchmark environment in the authors perspective. Safe-Control-Gym provides environments for three nonlinear dynamical systems: (i) Cart-Pole; (ii) 1D and 2D Quadrotor; (iii) 3D Quadrotor. The quadrotor environments have the following tasks:
@@
The environment uses PyBullet physics engine [136] to incorporate dynamic disturbances and external forces applied on the agent during simulation. These can be customized using the environment API to introduce necessary exogenous inputs and simulate uncertainty in the system. The environment API also allows for linear and quadratic custom constraints to be added to the states and control input. Overall, the environment provides all the necessary aspects to test and validate proposed safe RL methodologies.
@@ 5.2 Theoretical developments
Formal safety guarantees: One of the main advantages is the ability to provide formal safety guarantees. This provides a level of confidence that the agent will not deviate into unsafe states during learning or deployment. It is observed through this review that most model-based approaches provide strict guarantees compared to probabilistic or confidence bounds provided by model-free RL approaches. This is expected as the model-based RL has prior information about the dynamic structure of the system [53, 54, 67].
Risk mitigation during training: Lyapunov and barrier functions provide theoretical guarantees for the agent to actively avoid unsafe states or actions during the learning process as well. This proactive safety mechanism reduces system constraint violation (the risk of accidents or failures) during exploration while training and deployment, making them well-suited for applications in safety-critical domains [39, 41]. These approaches can be useful in deploying agents during the training process as well.
Adaptive safety and safe exploration: These methods can adapt to changes in the environment and system dynamics. Given most applications involve stochastic environments, the agent should be able adjust its behavior to accommodate variations and unforeseen events [104]. In [98], authors incorporate CBFs in multi-agent RL for cooperative and non-cooperative agent to achieve $0\%$ collision for a two man patrol problem. Safe exploration is a challenge in RL, especially in environments with high risks of failure leading to complete breakdown of the system. Lyapunov and barrier functions can guide exploration by constraining the agent’s actions within safe bounds. This ensures that an agent safely explores the environment by enforcing hard constraints (QP-based filters [105, 94]), providing formal guarantees during training and deployment [91, 96].
Smooth transitions (magnitude of control input) to safe states: Lyapunov barrier functions can enable smooth transitions of the agent from unsafe to safe states by tuning the accepted risk parameter and formulating a minimum control magnitude QP problem. This is particularly useful in scenarios where abrupt actions are required to avoid unsafe states could lead to instability or undesired actions [63, 91].
Figure 8: Safe Policy Optimization environment architecture [132]
@@ 5.3 Current Challenges
Conservatism and performance trade-off (exploration vs exploitation): RL primarily relies on exploring the environment to find an optimal (or effective) control policy. A key challenge in using Lyapunov or barrier functions in safe RL is the conservative exploration introduced by these safety constraints on the derived policy, often leading to sub optimal results. These constraints are designed to ensure forward invariance of the system within a safe set. This results in constrained or overly cautious/conservative exploration, limiting the agent’s ability to explore and learn efficiently. This creates a trade-off between safety and optimality creates a conundrum for the agent in achieving its objective. There have been several approaches to address this issue by perturbing the safe set [68, 83, 76]. However, most approaches present a tunable hyperparameter to adjust the allowable safe set perturbation. Formal or probabilistic bounds for the perturbation is an open area for research.
@@
Complexity and computational costs for systems with high dimensionality: Designing or approximating Lyapunov and/or barrier functions that accurately capture the safety requirements of a complex environment can be challenging since they require some prior information about the dynamic structure of the model. As complexity of the environment increases, finding appropriate Lyapunov or barrier functions becomes even more difficult. Furthermore, most approaches incorporate safety constraints as a constrained optimization problem to compute the safe action at each iteration which increases complexity as dimensionality of the problem increases [71, 97, 98, 83].
@@
Limited applicability for dynamic and unknown environments: Using Lyapunov and barrier functions in safe RL produces several challenges in case of partially observable environments. They often assume a certain level of structure or regularity in dynamics of the system, which might not hold in highly unpredictable or stochastic and partially observable environments. This is as open area of research given its applicability to real-world problems. This limitation is discussed as the drawback of using model-free RL approaches. How much data is sufficient to get an accurate approximation of CLF and CBF for the true system to provide formal guarantees, remains an open question. Furthermore, generalizing safety constraints to unknown and partially observable environments are scarcely present in literature. If the learned safety constraints are too tailored to the training environment, they might not transfer effectively to different contexts, leading to unsafe behavior when the agent encounters unknown scenarios.
@@
Current lack of real-world robustness for unmodeled dynamics and disturbances: While Lyapunov and barrier functions can provide guarantees in theory, real-world systems often have unmodeled dynamics, sensor noise or imperfect measurements, and external disturbances. These can lead to violations and often infeasible solutions to the assumed safety guarantees, resulting in unsafe behavior. Model-free approaches seem promising in this regard as shown in [75, 76, 104].
@@ 5.4 Future Work
Model-based approaches inherently seem suited to the incorporation of Lyapunov and barrier functions due to prior information available for the dynamic structure of the model. However, in most cases this is not possible due to incomplete model knowledge, unknown disturbances and information being only available while interacting with the environment. Model-free approaches present great opportunities in this regard. Theoretical gaps exist in terms of general conditions required for selecting the Lyapunov function as the value function [67, 68]. The general applicability of selecting a candidate Lyapunov function as value function remains to be formalized. Another area of research can be to quantify data requirements for sufficiently approximating the candidate Lyapunov or barrier functions. [67] provides an explanation to this problem by proposing use of Almost Lyapunov conditions, however, amount of data required or generalization error (performance in training compared to the entire state space) still remains an open area for research. Convergence guarantees while using deep or large NNs ($\geq 100$ hidden units) as function approximators is another promising area of research for RL algorithms. Using NNs of above-mentioned structure makes theoretical guarantees intractable. As observed in this review, several methods approximate the nonlinear dynamics to analytically obtain an expression for the Lie derivative employed in CLF and CBF QP formulations. Recent advances in approximating nonlinear dynamics as a linear system using Koopman theory [137] may be an interesting area to explore and help provide a generalized structure for this.
@@ 5.5 Potential and Current Applications
Lyapunov and barrier functions in safe RL have found significant application in robotics and autonomous systems, primarily ensuring collision avoidance, stability, and constraint adherence. Barrier functions integrated into RL have been effectively used for autonomous vehicles, providing formal guarantees for safety through collision avoidance, lane keeping, and speed control. For instance, CBFs guarantee safety-critical constraints, enabling vehicles to navigate complex traffic conditions safely while optimizing travel efficiency [38]. Lyapunov-based methods have been leveraged in multi-agent UAV systems for trajectory planning and formation control. They provide theoretical guarantees of system stability and collision avoidance, critical for complex multi-agent coordination tasks [3, 101]. Barrier functions have been used in industrial robots to ensure safety when interacting with dynamic environments or humans. These methods ensure physical safety without compromising operational efficiency, especially in collaborative robot settings. [72].
@@
In Cyber-Physical Systems (CPS), Lyapunov and barrier functions contribute significantly to maintaining system safety, stability, and reliability. Lyapunov functions have been successfully applied to maintain stability under load disturbances and mitigating risks due to cyber-attacks. Barrier functions protect against unsafe operating states, particularly during contingencies such as False Data Injection (FDI) attacks [17]. Barrier functions integrated with RL methods ensure safe control in chemical processes by maintaining operational constraints, such as pressure, temperature, and flow rates. Lyapunov stability techniques ensure convergence of RL-based controllers to safe equilibria [66].
@@
In medical applications, guaranteeing safety is paramount. Lyapunov and barrier functions ensure compliance with critical constraints and improve patient outcomes. Safe RL leveraging Lyapunov functions ensures stability and safety in automated drug delivery, for instance, insulin infusion in diabetic patients. It guarantees that RL-based controllers maintain blood glucose levels within a safe operational range, effectively preventing adverse events [138]. Barrier functions have been proposed to ensure surgical robots avoid hazardous regions or unsafe interactions with critical tissues, making robotic surgery safer and more reliable [139].
@@
Safe RL methods have critical applications in intelligent transportation systems, employing Lyapunov stability criterion has optimized urban traffic signals to reduce congestion while guaranteeing minimal waiting times and avoiding unsafe traffic configurations [140]. Barrier functions ensure safety-critical constraints such as speed limits and collision avoidance in railway networks. These applications minimize operational risks while optimizing performance and efficiency [141].
@@
Aerospace systems relies heavily on guaranteed safety and stability provided by Lyapunov and barrier function methodologies. Safe RL augmented with CLF have been employed to optimize satellite trajectory planning, ensuring orbital stability and collision-free maneuvers [142, 143]. Barrier functions in RL frameworks ensure adherence to critical safety constraints (e.g., angle-of-attack limits, airspeed criterion etc.) while maintaining optimal performance during aircraft operations [144].
@@
Most literature in this review focuses on use of safe RL in the robotics domain, this is primarily because mobile robotic systems require satisfying hard constraints and computing a stabilizing policy for general applications which cannot be guaranteed with vanilla RL. Furthermore, safe RL methods provide an opportunity for online learning in case of using QP based formulations, as discussed in [98, 96, 97]. A key domain that can benefit from safe RL with theoretical guarantees is power system control, which requires the RL algorithm to satisfy hard constraints for system states and dynamics [145]. There is limited in limited literature considering use of Lyapunov and barrier functions for safe RL in power systems. The most notable approach [146] considers use of barrier functions for emergency control during under voltage load shedding. Overall, nonlinear dynamical systems which require satisfying hard constraints for operation can greatly benefit from using safe RL.
@@ 6 Conclusion
Safe RL using Lyapunov and barrier functions work effectively in providing safety guarantees during the learning process and trained policy for agents. Use of Lyapunov and barrier functions in RL helps solve the key drawback of RL methods being implemented and widely adopted for safety critical applications. The review summarizes and presents insight to different approaches pertaining to use of: Lyapunov functions, barrier functions and combining the two approaches in RL. Most work in the field is fairly recent with model-free approaches being employed from 2017, which presents significant opportunities to bridge the gap between theory and simulations in RL with further applications to real-world systems.
7 Acknowledgments
The authors acknowledge use of ChatGPT to improve grammatical proficiency and formulation of grammatically correct sentences.

### v3

@@ 7 Discussion
Safe RL with Lyapunov or barrier functions is an effective and promising field that aims to ensure the safety of an agent while it learns a task through RL. Table 13 and 5 summarize the methodology of using Lyapunov and barrier functions in RL.
@@ 7.1 Simulation Benchmarks
Simulated benchmarks have proven to be a norm in implementing and validating data-driven ML and RL algorithms over the last few decades[dulac2021challenges]. The key challenges involved in creating a successful benchmark for these data-driven algorithms are reproducibility, ease of implementation, computational complexity (should be less to allow testing on all scales of hardware), incorporating stochastic behavior and sufficient environment complexity (indicating the environment goals should be sufficiently complex). We review some of the existing environments for benchmarking safe RL and address its usefulness in validating methodologies incorporating Lyapunov and barrier functions.
(a) (A) Island Navigation environment [leike2017aisafetygridworlds]
(b) (B) Constraints in Safe Gym [ray2019benchmarking]
@@
The earliest benchmark for safe RL approached the problem as a safe exploration towards a goal. DeepMind introduced AI Safety Gridworlds [leike2017aisafetygridworlds] and OpenAI released Safety Gym [ray2019benchmarking] as benchmarks for safe exploration problem in RL. Safety Gym proposes a constrained RL formulation for safe exploration and provide gym [brockman2016openai] environments [1606.01540] for high-dimensional continuous control in safe RL. Safety Gym was upgraded by introducing Safety-Gymnasium [ji2023safety] with increased functionality by using MuJoCo physics engine [todorov2012mujoco], higher customization, extended single and multi-agent scenarios. Safety-Gymnasium is built on Gymnasium [towers2024gymnasium] RL environment structure and allows for easy integration with RL libraries. Safe-Gymnasium also proposes SafePO which provides single and multi-agent approaches for constrained safe RL problems. The ideal benchmarking environment for reviewed methodologies should incorporate constraints, external disturbances, nonlinear dynamical models and low-level oriented tasks. For this reason Safe-Control-Gym [9849119] appears as the most suitable benchmark environment in the authors perspective. Safe-Control-Gym provides environments for three nonlinear dynamical systems: (i) Cart-Pole; (ii) 1D and 2D Quadrotor; (iii) 3D Quadrotor. The quadrotor environments have the following tasks:
@@
The environment uses PyBullet physics engine [coumans2019] to incorporate dynamic disturbances and external forces applied on the agent during simulation. These can be customized using the environment API to introduce necessary exogenous inputs and simulate uncertainty in the system. The environment API also allows for linear and quadratic custom constraints to be added to the states and control input. Overall, the environment provides all the necessary aspects to test and validate proposed safe RL methodologies.
@@ 7.2 Theoretical developments
Formal safety guarantees
One of the main advantages is the ability to provide formal safety guarantees. This provides a level of confidence that the agent will not deviate into unsafe states during learning or deployment. It is observed through this review that most model-based approaches provide strict guarantees compared to probabilistic or confidence bounds provided by model-free RL approaches. This is expected as the model-based RL has prior information about the dynamic structure of the system [vamvoudakis2015asymptotically, kamalapurkar2016model, chang2021stabilizing].
Risk mitigation during training
Lyapunov and barrier functions provide theoretical guarantees for the agent to actively avoid unsafe states or actions during the learning process as well. This proactive safety mechanism reduces system constraint violation (the risk of accidents or failures) during exploration while training and deployment, making them well-suited for applications in safety-critical domains [garcia2015comprehensive, brunke2022safe]. These approaches can be useful in deploying agents during the training process as well.
Adaptive safety and safe exploration
These methods can adapt to changes in the environment and system dynamics. Given most applications involve stochastic environments, the agent should be able adjust its behavior to accommodate variations and unforeseen events [cheng2023safe]. In [cai2021safe], authors incorporate CBFs in multi-agent RL for cooperative and non-cooperative agent to achieve $0\%$ collision for a two man patrol problem. Safe exploration is a challenge in RL, especially in environments with high risks of failure leading to complete breakdown of the system. Lyapunov and barrier functions can guide exploration by constraining the agent’s actions within safe bounds. This ensures that an agent safely explores the environment by enforcing hard constraints (QP-based filters [hou2024safe, liu2025safe]), providing formal guarantees during training and deployment [marvi2021safe, emam2021safe].
Smooth transitions (magnitude of control input) to safe states
Lyapunov barrier functions can enable smooth transitions of the agent from unsafe to safe states by tuning the accepted risk parameter and formulating a minimum control magnitude QP problem. This is particularly useful in scenarios where abrupt actions are required to avoid unsafe states could lead to instability or undesired actions [10155918, marvi2021safe].
Figure 8: Safe Policy Optimization environment architecture [ji2023safety]
@@ 7.3 Current Challenges
RL primarily relies on exploring the environment to find an optimal (or effective) control policy. A key challenge in using Lyapunov or barrier functions in safe RL is the conservative exploration introduced by these safety constraints on the derived policy, often leading to sub optimal results. These constraints are designed to ensure forward invariance of the system within a safe set. This results in constrained or overly cautious/conservative exploration, limiting the agent’s ability to explore and learn efficiently. This creates a trade-off between safety and optimality resulting in a conundrum for the agent in achieving its objective safely. There have been several approaches to address this issue by perturbing the safe set [westenbroek2022lyapunov, ohnishi2019barrier, chow2019lyapunov]. However, most approaches present a tunable hyperparameter to adjust the allowable safe set perturbation. Formal or probabilistic bounds for the magnitude of perturbation for the safe set is an open area for research.
Complexity and computational costs for systems with high dimensionality
@@
Designing or approximating Lyapunov and/or barrier functions that accurately capture the safety requirements of a complex environment can be challenging since they require some prior information about the dynamic structure of the model. As complexity of the environment increases, finding appropriate Lyapunov or barrier functions becomes even more difficult. Furthermore, most approaches incorporate safety constraints as a constrained optimization problem to compute the safe action at each iteration which increases complexity as dimensionality of the problem increases [westenbroek2020learning, ma2021model, cai2021safe, ohnishi2019barrier].
Limited applicability for dynamic and unknown environments
@@
Using Lyapunov and barrier functions in safe RL produces several challenges in case of partially observable environments. They often assume a certain level of structure or regularity in dynamics of the system, which might not hold in highly unpredictable or stochastic and partially observable environments. This is as open area of research given its applicability to real-world problems. This limitation is discussed as the drawback of using model-free RL approaches. How much data is sufficient to get an accurate approximation of CLF and CBF for the true system to provide formal guarantees, remains an open question. Furthermore, generalizing safety constraints to unknown and partially observable environments are scarcely present in literature. If the learned safety constraints are too tailored to the training environment, they might not transfer effectively to different contexts, leading to unsafe behavior when the agent encounters unknown scenarios.
Current lack of real-world robustness for unmodeled dynamics and disturbances
@@
While Lyapunov and barrier functions can provide guarantees in theory, real-world systems often have unmodeled dynamics, sensor noise or imperfect measurements, and external disturbances. These can lead to violations and often infeasible solutions to the assumed safety guarantees, resulting in unsafe behavior. Model-free approaches seem promising in this regard as shown in [chow2018lyapunov, chow2019lyapunov, cheng2023safe].
@@ 7.4 Future Work
Model-based approaches inherently seem suited to the incorporation of Lyapunov and barrier functions due to prior information available for the dynamic structure of the model. However, in most cases this is not possible due to incomplete model knowledge, unknown disturbances and information being only available while interacting with the environment. Model-free approaches present great opportunities in this regard. Theoretical gaps exist in terms of general conditions required for selecting the Lyapunov function as the value function [chang2021stabilizing, westenbroek2022lyapunov]. The general applicability of selecting a candidate Lyapunov function as value function remains to be formalized. Another area of research can be to quantify data requirements for sufficiently approximating the candidate Lyapunov or barrier functions. [chang2021stabilizing] provides an explanation to this problem by proposing use of Almost Lyapunov conditions, however, amount of data required or generalization error (performance in training compared to the entire state space) still remains an open area for research. Convergence guarantees while using deep or large NNs ($\geq 100$ hidden units) as function approximators is another promising area of research for RL algorithms. Using NNs of above-mentioned structure makes theoretical guarantees intractable. As observed in this review, several methods approximate the nonlinear dynamics to analytically obtain an expression for the Lie derivative employed in CLF and CBF QP formulations. Recent advances in approximating nonlinear dynamics as a linear system using Koopman theory [brunton2021modern] may be an interesting area to explore and help provide a generalized structure for this.
@@ 7.5 Potential and Current Applications
Lyapunov and barrier functions in safe RL have found significant application in robotics and autonomous systems, primarily ensuring collision avoidance, stability, and constraint adherence. Barrier functions integrated into RL have been effectively used for autonomous vehicles, providing formal guarantees for safety through collision avoidance, lane keeping, and speed control. For instance, CBFs guarantee safety-critical constraints, enabling vehicles to navigate complex traffic conditions safely while optimizing travel efficiency [ames2019control]. Lyapunov-based methods have been leveraged in multi-agent UAV systems for trajectory planning and formation control. They provide theoretical guarantees of system stability and collision avoidance, critical for complex multi-agent coordination tasks [wang2022distributed, wang2022ensuring]. Barrier functions have been used in industrial robots to ensure safety when interacting with dynamic environments or humans. These methods ensure physical safety without compromising operational efficiency, especially in collaborative robot settings [long2025certifying].
@@
In Cyber-Physical Systems (CPS), Lyapunov and barrier functions contribute significantly to maintaining system safety, stability, and reliability. Lyapunov functions have been successfully applied to maintain stability under load disturbances and mitigating risks due to cyber-attacks. Barrier functions protect against unsafe operating states, particularly during contingencies such as False Data Injection (FDI) attacks [10252253]. Barrier functions integrated with RL methods ensure safe control in chemical processes by maintaining operational constraints, such as pressure, temperature, and flow rates. Lyapunov stability techniques ensure convergence of RL-based controllers to safe equilibrium [yao2025lyapunov].
@@
In medical applications, guaranteeing safety is paramount. Lyapunov and barrier functions ensure compliance with critical constraints and improve patient outcomes. Safe RL leveraging Lyapunov functions ensures stability and safety in automated drug delivery, for instance, insulin infusion in diabetic patients. It guarantees that RL-based controllers maintain blood glucose levels within a safe operational range, effectively preventing adverse events [yu2025deep]. Barrier functions have been proposed to ensure surgical robots avoid hazardous regions or unsafe interactions with critical tissues, making robotic surgery safer and more reliable [fan2024learn].
@@
Safe RL methods have critical applications in intelligent transportation systems, employing Lyapunov stability criterion has optimized urban traffic signals to reduce congestion while guaranteeing minimal waiting times and avoiding unsafe traffic configurations [wei2021recent]. Barrier functions ensure safety-critical constraints such as speed limits and collision avoidance in railway networks. These applications minimize operational risks while optimizing performance and efficiency [lovetei2022environment].
@@
Aerospace systems relies heavily on guaranteed safety and stability provided by Lyapunov and barrier function methodologies. Safe RL augmented with CLF have been employed to optimize satellite trajectory planning, ensuring orbital stability and collision-free maneuvers [yu2024safe, yi2025free]. Barrier functions in RL frameworks ensure adherence to critical safety constraints (e.g., angle-of-attack limits, airspeed criterion etc.) while maintaining optimal performance during aircraft operations [jiang2023safely].
@@
Most literature in this review focuses on use of safe RL in the robotics domain, this is primarily because mobile robotic systems require satisfying hard constraints and computing a stabilizing policy for general applications which cannot be guaranteed with vanilla RL. Furthermore, safe RL methods provide an opportunity for online learning in case of using QP based formulations, as discussed in [cai2021safe, emam2021safe, ma2021model]. A key domain that can benefit from safe RL with theoretical guarantees is power system control, which requires the RL algorithm to satisfy hard constraints for system states and dynamics [chen2022reinforcement]. There is limited in limited literature considering use of Lyapunov and barrier functions for safe RL in power systems. The most notable approach [vu2021barrier] considers use of barrier functions for emergency control during under voltage load shedding. Overall, nonlinear dynamical systems which require satisfying hard constraints for operation can greatly benefit from using safe RL.
@@ 8 Conclusion
Safe RL using Lyapunov and barrier functions work effectively in providing safety guarantees during the learning process and deployment for agents. Use of Lyapunov and barrier functions in RL helps solve the key drawback (safety and stabilizing policy) of RL methods being implemented and widely adopted for safety critical applications. The review summarizes and presents insight to different approaches pertaining to use of: Lyapunov functions, barrier functions and combining the two approaches in RL. Most work in the field is fairly recent with model-free approaches being employed from 2017, which presents significant opportunities to bridge the gap between theory and simulations in RL with further applications to real-world systems.